A department store held a sale to sell all of the 214 winter jackets that remained after the season ended. Until noon, each jacket in the store was priced at $31.95. At noon, the price of the jackets was further reduced to $18.95. After the last jacket was sold, the total receipts for the clearance sale were $5108.30. How many jackets were sold before noon and how many were sold after noon?
Answer by checkley71(8403)

x+y=214 x being the sales before noon & y those after noon. or-
y=214-x
31.95x+18.95(214-x)=5108.30
31.95x+4055.30-18.95x=5108.30
13x=5108.30-4055.30
13x=1053
x=1053/13
x=81 jackets sold befor noon.
214-81=133 jackets sold after noon.
proof
31.95*81+18.95*133=5108.30
2587.95+2520.35=5108.30
5108.30=5108.30